# Biased random walk on dynamical percolation

### v2

@@ 4.3 Asymptotic derivative of the speed
We start both walks from 00 and we let them both attempt jumps at the points of a Poisson process $\mathcal{P}=(\mathcal{P}_{t})_{t\geq 0}$ of rate 1. We also let both environments evolve together until the first very bad point defined below and afterwards we couple the environments by using the same rate $\mu$ Poisson process for the removal of copies of edges of their respective infected sets. Whenever a clock of $\mathcal{P}$ rings indicating the jump attempt at time $t$ of both walkers, we sample a random variable $U$ uniformly on $[0,1]$ and proceed as follows:
@@
$\begin{split}\lim_{\varepsilon\to 0}\frac{\mathbb{P}\!\left(A_{\varepsilon}\right)}{\varepsilon}&=\mathbb{E}\!\left[\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\left(1-\Big{(}1-\varepsilon(2d-2)\exp(-\lambda)+\mathcal{O}(\varepsilon e^{-2\lambda}+\varepsilon^{2})\Big{)}^{\mathcal{U}_{a}(\tau_{1})}\right)\right]\\ &=(2d-2)e^{-\lambda}\cdot\mathbb{E}\!\left[\mathcal{U}_{a}(\tau_{1})\right]+\mathcal{O}(e^{-2\lambda}),\end{split}$ (4.19)
@@ 5 Strict monotonicity of the speed for large $\mu$ or $p$ close to $1$
Recall the function $f(\lambda)$ from (3) as well as $Z_{\lambda}$ from (1.1) and $Z^{\prime}_{\lambda}=e^{\lambda}-e^{-\lambda}$. Moreover, for $k,\ell,m\in\mathbb{N}$, we write
$f_{k,\ell,m}(\lambda):=(k-\ell)e^{\lambda(k-\ell)}\left(\frac{2d}{Z_{\lambda}}\right)^{m}\left(k-\ell-m\cdot\frac{Z_{\lambda}^{\prime}}{Z_{\lambda}}\right)\mathbb{P}_{0}\big{(}(R,L)=(k,\ell),\mathcal{U}(\tau_{1})=m\big{)}$
@@
$f^{\prime}(\lambda)=\sum_{m\in\mathbb{N}}\sum_{k+\ell\leq m}f_{k,\ell,m}(\lambda).$
@@
It suffices to prove that $f^{\prime}(\lambda)$ is strictly positive for all $\lambda>0$ provided $\mu$ is sufficiently large. For all $i,j\in\mathbb{N}$ and $m\geq 2$ using Lemma 2.3 we have
$\mathbb{P}_{0}\big{(}(R,L)=(i,j),\mathcal{U}(\tau_{1})=m)\leq\exp(-c_{\mu}m\big{)}$ (5.1)
@@
$\mathbb{P}_{0}\big{(}(R,L)=(1,0),\mathcal{U}(\tau_{1})=1\big{)}=\mathbb{P}_{0}\big{(}(R,L)=(0,1),\mathcal{U}(\tau_{1})=1\big{)}\geq p\cdot\frac{1}{2d}\cdot\frac{\mu}{\mu+1}\geq\frac{p}{4d}\,,$ (5.2)
@@
$A_{m}:=\big{\{}(k,\ell)\in\mathbb{N}^{2}:k+\ell\leq m\ \text{ and }\ k-\ell\leq m-1\big{\}}.$
@@
$\sum_{m\geq 2}\sum_{(k,\ell)\in A_{m}}|f_{k,\ell,m}(\lambda)|\leq\sum_{m\geq 2}2m^{3}\cdot e^{\lambda(m-1)}\left(\frac{2d}{Z_{\lambda}}\right)^{m}e^{-\textup{c}_{\mu}m}\leq e^{-\lambda}\sum_{m\geq 2}2m^{3}(2d)^{m}e^{-\textup{c}_{\mu}m}.$
Since $c_{\mu}\to\infty$ as $\mu\to\infty$, by taking $\mu$ sufficiently large we can make the sum above as small as desired. Using (5.2) and that $d\geq 2$ we now notice that there exists a positive constant $c$ such that for all $\lambda>0$
$f_{1,0,1}(\lambda)+f_{0,1,1}(\lambda)\geq ce^{-\lambda}.$
We thus deduce that for $\mu$ sufficiently large we get for all $\lambda>0$
$\sum_{m\geq 2}\sum_{(k,\ell)\in A_{m}}|f_{k,\ell,m}(\lambda)|\leq\frac{1}{2}(f_{1,0,1}(\lambda)+f_{0,1,1}(\lambda)).$ (5.3)
Taking now into account the summands not contained in $A_{m}$, since $f_{m,0,m}(\lambda)\geq f_{0,m,m}(\lambda)$ for all $\lambda>0$ and $m\in\mathbb{N}$ and
$f_{0,m,m}(\lambda)=m^{2}e^{-\lambda m}\left(\frac{2d}{Z_{\lambda}}\right)^{m}\left(1+\frac{Z_{\lambda}^{\prime}}{Z_{\lambda}}\right)\mathbb{P}_{0}\big{(}(R,L)=(m,0),\mathcal{U}(\tau_{1})=m\big{)}\geq 0,$ (5.4)
@@
$f^{\prime}(\lambda)\geq\frac{1}{2}(f_{1,0,1}(\lambda)+f_{0,1,1}(\lambda)).$
@@
$f_{1,0,1}(\lambda)+f_{0,1,1}(\lambda)\geq\frac{2d}{Z_{\lambda}}\cdot\frac{(2d-2)(e^{\lambda}+e^{-\lambda})+4}{Z_{\lambda}}\cdot\frac{p}{4d}>0$
@@
uniformly for all $\lambda\in(0,\lambda_{0}]$. Recall the expression for $v^{\prime}(\lambda,p)$ from Lemma 3.4. We want to compare $\mathbb{E}_{\lambda,p}\!\left[(X_{\tau_{1}}^{1})^{2}\right]$ to $\mathbb{E}_{\lambda,1}\!\left[(X_{\tau_{1}}^{1})^{2}\right]$ and also $\mathbb{E}_{\lambda,p}\!\left[X_{\tau_{1}}^{1}\cdot\mathcal{U}(\tau_{1})\right]$ to $\mathbb{E}_{\lambda,1}\!\left[X_{\tau_{1}}^{1}\cdot\mathcal{U}(\tau_{1})\right]$. To do this, we couple the walks in the two environments by letting them attempt the same jumps at the same times and using the same Poisson process of rate $\mu$ to remove copies of edges from their infected sets. We let $\tau_{1}$ be their first regeneration time. We now let $\kappa$ be the first time when the walk in the $p$-dynamical percolation process attempts a jump along a closed edge. Then up until time $\kappa$ the two walks are in the same location. Note that $\kappa$ stochastically dominates a geometric random variable of parameter $1-p$, because for all $s<t$ and all edges $e$ we have
@@
By a union bound we get
$\mathbb{P}\!\left(\kappa<\mathcal{U}(\tau_{1})\right)\leq\mathbb{P}\!\left(\mathcal{U}(\tau_{1})>\frac{1}{\sqrt{1-p}}\right)+\mathbb{P}\!\left(\kappa<\frac{1}{\sqrt{1-p}}\right)\leq C_{\mu}\cdot\sqrt{1-p},$
where $C_{\mu}$ is a constant depending on $\mu$. Using this and the bound $|X_{\tau_{1}}^{1}|\leq\mathcal{U}(\tau_{1})$ we get
$|\mathbb{E}_{\lambda,p}\!\left[(X_{\tau_{1}}^{1})^{2}\right]-\mathbb{E}_{\lambda,1}\!\left[(X_{\tau_{1}}^{1})^{2}\right]|\leq 2\mathbb{E}_{\lambda,p}\!\left[(\mathcal{U}(\tau_{1}))^{2}{\text{\Large$\mathfrak{1}$}}(\kappa<\mathcal{U}(\tau_{1}))\right].$
By the Cauchy-Schwarz inequality and the exponential tails of $\mathcal{U}(\tau_{1})$ uniformly in $p$ by Lemma 2.3 (using the obvious bound $\mathcal{U}(\tau_{1})\leq\mathcal{U}_{a}(\tau_{1})$) we deduce
$\mathbb{E}_{\lambda,p}\!\left[(\mathcal{U}(\tau_{1}))^{2}{\text{\Large$\mathfrak{1}$}}(\kappa<\mathcal{U}(\tau_{1}))\right]\leq C_{\mu}^{\prime}(1-p)^{1/4}.$
@@
[Acknowledgments] We thank Frank den Hollander and Remco van der Hofstad for valuable discussions. We also thank the anonymous referees for valuable comments and suggestions. {funding} DS acknowledges the DAAD PRIME program for financial support.

### v3

@@ 4.3 Asymptotic derivative of the speed
We start both walks from 00 and we let them both attempt jumps at the points of a Poisson process $\mathcal{P}=(\mathcal{P}_{t})_{t\geq 0}$ of rate 1. We also let both environments evolve together until the first very bad point defined below and afterwards we couple the environments by using the same rate $\mu$ Poisson process for the removal of copies of edges of their respective infected sets.
Whenever a clock of $\mathcal{P}$ rings indicating the jump attempt at time $t$ of both walkers, we sample a random variable $U$ uniformly on $[0,1]$ and proceed as follows:
@@
$\begin{split}\lim_{\varepsilon\to 0}\frac{\mathbb{P}\!\left(A_{\varepsilon}\right)}{\varepsilon}&=\mathbb{E}\!\left[\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\left(1-\Big{(}1-\varepsilon(2d-2)\exp(-\lambda)+\mathcal{O}(\varepsilon e^{-2\lambda}+\varepsilon^{2}e^{-\lambda})\Big{)}^{\mathcal{U}_{a}(\tau_{1})}\right)\right]\\ &=(2d-2)e^{-\lambda}\cdot\mathbb{E}\!\left[\mathcal{U}_{a}(\tau_{1})\right]+\mathcal{O}(e^{-2\lambda}),\end{split}$ (4.17)
@@ 5 Strict monotonicity of the speed for large $\mu$ or $p$ close to $1$
Recall the function $f(\lambda)$ from (3) as well as $Z_{\lambda}$ from (1.1) and $Z^{\prime}_{\lambda}=e^{\lambda}-e^{-\lambda}$, and $p_{0}$ from (3.15). For $k_{a},\ell_{a},k,\ell\in\mathbb{N}_{0}$ and $m\in\mathbb{N}$, we write
$f_{k_{a},\ell_{a},k,\ell,m}(\lambda):=(k-\ell)e^{\lambda(k_{a}-\ell_{a})}\left(\frac{2d}{Z_{\lambda}}\right)^{m}\left(k_{a}-\ell_{a}-m\cdot\frac{Z_{\lambda}^{\prime}}{Z_{\lambda}}\right)p_{0}(k_{a},\ell_{a},k,\ell,m)$ (5.1)
@@
$f^{\prime}(\lambda)=\sum_{m\in\mathbb{N}}\sum_{\begin{subarray}{c}k_{a}+\ell_{a}\leq m\\ k\leq k_{a},\ell\leq\ell_{a}\end{subarray}}f_{k_{a},\ell_{a},k,\ell,m}(\lambda).$
@@
It suffices to prove that $f^{\prime}(\lambda)$ is strictly positive for all $\lambda>0$ provided $\mu$ is sufficiently large. For all $k_{a},\ell_{a},k,\ell\in\mathbb{N}_{0}$ and $m\geq 2$ using Lemma 2.3 we have
$p_{0}(k_{a},\ell_{a},k,\ell,m)\leq\exp(-c_{\mu}m\big{)}$ (5.2)
@@
$p_{0}(1,0,1,0,1)=p_{0}(0,1,0,1,1)\geq p\cdot\frac{1}{2d}\cdot\frac{\mu}{\mu+1}\geq\frac{p}{4d}\,,$ (5.3)
@@
$A_{m}:=$$\big{\{}(k_{a},\ell_{a},k,\ell)\in\mathbb{N}_{0}^{4}:k\leq k_{a}\text{ and }\ell\leq\ell_{a}\text{ and }k_{a}+\ell_{a}\leq m\big{\}}$
$\setminus\{(m,0,k,0),(0,m,0,\ell):k,\ell\leq m\}.$
@@
$\sum_{m\geq 2}\sum_{(k_{a},\ell_{a},k,\ell)\in A_{m}}|f_{k_{a},\ell_{a},k,\ell,m}(\lambda)|$ $\leq\sum_{m\geq 2}2m^{6}\cdot e^{\lambda(m-1)}\left(\frac{2d}{Z_{\lambda}}\right)^{m}e^{-\textup{c}_{\mu}m}$
$\leq e^{-\lambda}\sum_{m\geq 2}2m^{6}(2d)^{m}e^{-\textup{c}_{\mu}m}.$
Since $c_{\mu}\to\infty$ as $\mu\to\infty$, by taking $\mu$ sufficiently large we can make the sum above as small as desired. Using (5.3) and that $d\geq 2$ we claim that there exists a positive constant $c$ such that for all $\lambda>0$
$f_{1,0,1,0,1}(\lambda)+f_{0,1,0,1,1}(\lambda)\geq ce^{-\lambda}.$ (5.4)
To see this, note that for $\lambda\rightarrow 0$, the left-hand side in (5.4) converges to $02p_{0}(1,0,1,0,1)>$ 0. For $\lambda\rightarrow\infty$, the estimate in (5.4) follows from (5.1). This allows to conclude (5.4) using continuity in $\lambda$. We thus deduce that for $\mu$ sufficiently large we get for all $\lambda>0$
$\sum_{m\geq 2}\sum_{(k_{a},\ell_{a},k,\ell)\in A_{m}}|f_{k_{a},\ell_{a},k,\ell,m}(\lambda)|\leq\frac{1}{2}(f_{1,0,1,0,1}(\lambda)+f_{0,1,0,1,1}(\lambda)).$ (5.5)
Taking now into account the summands not contained in $A_{m}$, note that
$f_{m,0,n,0,m}(\lambda)\geq f_{0,m,0,n,m}(\lambda)$ (5.6)
for all $\lambda>0$ and $m\in\mathbb{N}$ with $n\leq m$. Moreover,
$f_{0,m,0,n,m}(\lambda)=nme^{-\lambda m}\left(\frac{2d}{Z_{\lambda}}\right)^{m}\left(1+\frac{Z_{\lambda}^{\prime}}{Z_{\lambda}}\right)p_{0}(0,m,0,n,m)\geq 0.$ (5.7)
@@
$f^{\prime}(\lambda)\geq\frac{1}{2}(f_{1,0,1,0,1}(\lambda)+f_{0,1,0,1,1}(\lambda)).$
@@
$f_{1,0,1,0,1}(\lambda)+f_{0,1,0,1,1}(\lambda)\geq\frac{2d}{Z_{\lambda}}\cdot\frac{(2d-2)(e^{\lambda}+e^{-\lambda})+4}{Z_{\lambda}}\cdot\frac{p}{4d}>0$
@@
uniformly for all $\lambda\in(0,\lambda_{0}]$. Recall the expression for $v^{\prime}(\lambda,p)$ from Lemma 3.5. We want to compare $\mathbb{E}_{\lambda,p}\!\left[X_{\tau_{1}}^{1}(R_{a}-L_{a})\right]$ to $\mathbb{E}_{\lambda,1}\!\left[X_{\tau_{1}}^{1}(R_{a}-L_{a})\right]$ and also $\mathbb{E}_{\lambda,p}\!\left[X_{\tau_{1}}^{1}\cdot\mathcal{U}_{a}(\tau_{1})\right]$ to $\mathbb{E}_{\lambda,1}\!\left[X_{\tau_{1}}^{1}\cdot\mathcal{U}_{a}(\tau_{1})\right]$. To do this, we couple the walks in the two environments by letting them attempt the same jumps at the times $(T_{i})$ of a Poisson process of rate $1$ and using the same Poisson process of rate $\mu$ to remove copies of edges from their infected sets. We let $\tau_{1}$ be their first regeneration time. We now let $\kappa$ be the index of the first jump time when the walk in the $p$-dynamical percolation process attempts a jump along a closed edge. Then up until time $T_{\kappa}$ the two walks are in the same location. Note that $\kappa$ stochastically dominates a geometric random variable of parameter $1-p$, because for all $s<t$ and all edges $e$ we have
@@
Writing $\mathcal{U}_{a}(t)$ for the number of attempted jumps in the interval $[0,t]$, by a union bound we get
$\mathbb{P}\!\left(\kappa<\mathcal{U}_{a}(\tau_{1})\right)\leq\mathbb{P}\!\left(\mathcal{U}_{a}(\tau_{1})>\frac{1}{\sqrt{1-p}}\right)+\mathbb{P}\!\left(\kappa<\frac{1}{\sqrt{1-p}}\right)\leq C_{\mu}\cdot\sqrt{1-p},$
where $C_{\mu}$ is a constant depending on $\mu$. Using this and the bound $|X_{\tau_{1}}^{1}|\leq\mathcal{U}_{a}(\tau_{1})$ we get
$|\mathbb{E}_{\lambda,p}\!\left[X_{\tau_{1}}^{1}(R_{a}-L_{a})\right]-\mathbb{E}_{\lambda,1}\!\left[X_{\tau_{1}}^{1}(R_{a}-L_{a})\right]|\leq 2\mathbb{E}_{\lambda,p}\!\left[(\mathcal{U}_{a}(\tau_{1}))^{2}{\text{\Large$\mathfrak{1}$}}(\kappa<\mathcal{U}_{a}(\tau_{1}))\right].$
By the Cauchy-Schwarz inequality and the exponential tails of $\mathcal{U}_{a}(\tau_{1})$ uniformly in $p$ by Lemma 2.3, we deduce
$\mathbb{E}_{\lambda,p}\!\left[(\mathcal{U}_{a}(\tau_{1}))^{2}{\text{\Large$\mathfrak{1}$}}(\kappa<\mathcal{U}_{a}(\tau_{1}))\right]\leq C_{\mu}^{\prime}(1-p)^{1/4}.$
@@
[Acknowledgments] We thank Frank den Hollander and Remco van der Hofstad for valuable discussions. We also thank the anonymous referees for valuable comments and suggestions. Moreover, we are grateful to Jan Nagel for pointing out a mistake in an earlier version. {funding} DS acknowledges the DAAD PRIME program for financial support.